cos(sin^−1(x) − tan^−1(y))
Answer by stanbon(75887)

cos(sin^−1(x) − tan^−1(y))
-----
cos(sin^-1(x))
Let sin(t) = x
Then : cos(t) = sqrt(1-sin^2(t)) = sqrt(1-x^2)
----------
cos(tan^-1(y))
Let tan(t) = y/1
Then : cos(t) = 1/sqrt(y^2+1)
==========
Ans cos(sin^−1(x) − tan^−1(y)) = [sqrt(1-x^2)-[1/sqrt(y^2+1)]]
-----------
Cheers,
Stan H.
-----------